# Mixture of Cluster-conditional LoRA Experts for Vision-language Instruction Tuning

### 08 Feb 2024 v3

@@ Abstract
Instruction tuning of the Large Vision-language Models (LVLMs) has revolutionized the development of versatile models with zero-shot generalization across a wide range of downstream vision-language tasks. However, diversity of training tasks of different sources and formats would lead to inevitable task conflicts, where different tasks conflicts for the same set of model parameters, resulting in sub-optimal instruction-following abilities. To address that, we propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel Mixture of Experts (MoE) architecture designed to activate the task-customized model parameters based on the instruction clusters. A separate universal expert is further incorporated to improve the generalization capabilities of MoCLE for novel instructions. Extensive experiments on 12 zero-shot tasks demonstrate the effectiveness of MoCLE.
$*{}^{*}$ Equal contribution. ${}^{\dagger}$ Corresponding author.
@@ 1 Introduction
There has been a continuously increasing trend to develop intelligent assistants that can follow human instructions [4, 42, 8], and instruction tuning [55] has become a successful paradigm leveraging large-scale well-formatted instruction data to empower Large Language Models (LLMs) to follow human instructions, demonstrating strong generalization for novel unseen tasks [36]. Similarly, efforts have been made to introduce similar capabilities to Large Vision-language Models (LVLMs), including LLaVA [33], MiniGPT-4 [67] and InstructBLIP [11], which connects a trainable visual encoder with a frozen LLM and endows the LVLMs with the instruction-following ability by fine-tuning on multi-modal instruction datasets.
@@
The ability of generalizing to novel unseen instructions necessitates multi-task instruction tuning, i.e., training on a diverse collection of instruction-following tasks, which has been observed in both the LLMs [45, 53, 10] and LVLMs [3, 63, 28]. However, complexity of various instruction tasks brings difficulties for the model fine-tuning. Specifically, Wei et al. [55] observe that under certain model sizes, multi-task instruction tuning even fails to bring performance gains for zero-shot tasks compared to the original models, mainly attributed to the negative transfer phenomenon [62] during multi-task instruction tuning, where the model struggles to optimize the losses of multiple conflicted tasks, leading to sub-optimal performance for these tasks.
@@
Figure 1: Performance of the instruction-finetuned LVLMs on zero-shot tasks, where larger values indicate better performance. Only 3 out of 8 tasks benefit from instruction tuning from all the data, while the task experts show better performance on the other 5 tasks (i.e., Flickr 30K, HM, ScienceQA, IconQA, and VizWiz).
@@
Similarly, tasks for vision-language instruction tuning (e.g., visual question answering and image captioning) focus on different perspectives of LVLMs, and thus, resulting in conflicts when sharing all the parameters as most studies adopt. In our preliminary studies shown in Figure 1, we first split the instruction data into two disjoint subsets, including “cap” for image captioning and “vqa” for visual question answering. We then train InstructBLIP [11] using LoRA [18] on 3 kinds of data: “cap”, “vqa” and “full” (the full data) separately, to obtain 3 set of parameters (i.e., task experts). Following the held-out evaluation protocol in [11], we evaluate these experts on the unseen datasets/tasks with the best expert. As shown in Figure 1, on 5 out of the 8 downstream tasks, InstructBLIP instruction-tuned on all the data is outperformed by the task expert that is finetuned on only a subset of data. Among the 5 tasks, Flickr30k belongs to “cap”, and SciQA, IconQA and VizWiz belong to “vqa”, revealing that instruction tuning on the similar tasks brings positive transfer to downstream tasks, while training on the full data with dissimilar tasks could hurt generalization performance.
@@
The task experts above is a naive solution to negative transfer, where we manually partition the training tasks and train each expert individually. However, there are several limitations of this strategy, (1) The taxonomies such as “vqa” and “cap” require human expertise and become difficult to scale as the number of tasks grows; (2) The ability to generalize to unseen tasks is inhibited, as we do not know which expert to choose for novel tasks and some new tasks might benefit from multiple training tasks (e.g., VSR and TextVQA as in Figure 1). In this regard, specialization and generalization of LMMs becomes a dilemma.
@@
Our work aims to develop an automatic and practical partition strategy as well as a network architecture that can strike a balance between specialization and generalization. We propose Mixture of Cluster-conditional LoRA Experts (MoCLE) for vision-language instruction tuning. Under the proposed MoCLE framework, we first cluster instructions of all the training data into several centroids via a pretrained clustering model. In this way, similar tasks that could bring positive transfer to each other are grouped together in a cluster (more justification for the use of instruction clusters is detailed in Sec. 3.3). Then using the cluster as a condition, at each layer of the model, a router dispatches the input data to one of the specialized task experts and also a universal expert that is shared among all the data. Finally, these two set of outputs are combined and passed to the next layer. As we activate a specialized expert for a group of similar tasks, tasks that are less alike are learned via separate experts, mitigating task conflicts. Meanwhile, since the universal expert trained on all tasks also contributes to the model outputs, we can enjoy generalization and specialization simultaneously.
@@
We validate the effectiveness of MoCLE on the state-of-the-art InstructBLIP [11]. MoCLE achieves remarkable improvement on the held-out tasks [11] including image captioning (e.g., Flickr30K [58]), text-based visual question answering (e.g., TextVQA [50]), knowledge-based visual question answering (e.g., ScienceQA [38]) and novel visual reasoning (e.g., Visual Spatial Reasoning [32]).
@@
    We identify the negative transfer phenomenon [34, 65] of task conflicts during instruction tuning of LVLMs.
@@
    We propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel parameter-efficient finetuning framework suitable for the vision-language instruction tuning, to mitigate task conflicts and enjoy the benefits of huge data training simultaneously.
@@
    The proposed MoCLE method achieves remarkable performance gains on the unseen held-out tasks compared with a strong baseline model InstructBLIP [11].
@@ 2 Related Work
Multi-task Instruction Tuning.
@@
Instruction tuning, as elucidated in [45, 55], is a technique where a language model is finetuned over a collection of tasks and instruction templates that tells the model the intention of the task. The primary goal of this technique is to teach the model to understand the relationship between the instruction and the input/output pairs, so that the model could generalize to unseen task given new instructions. Increasing the number of diverse instructions [45], the number of tasks [53, 10], and diversity of data [66] have been shown to improve the performance. However, Wei et al. [55] observed under certain model sizes, instruction tuning of language models fail to bring performance gains to unseen tasks when compared with untuned base models. They argue this is due to the fact all model capacity is used to learn the mixture of instruction tuning tasks. In this work, we aim to mitigate such issue in vision-language instruction tuning by specialized experts.
Large Multi-modal Models.
@@
To inherit strong reasoning and generalization abilities of LLMs and meanwhile understand visual inputs, recent works including Flamingo [2], BLIP-2 [26], LLaMA-adapter [61] connect LLMs to the visual modality with a frozen visual encoder and a trainable module, which is then finetuned with plain image-text pairs. To further improve instruction following abilities and response quality, LLaVA finetunes the whole LLM on vision-language conversational data generated by GPT-4. MiniGPT-4 performs training on ChatGPT-generated image captions instead of traditional caption data that are much shorter. InstructBLIP collects a wide range of academic VQA datasets and manually crafts 10 to 15 distinct instruction templates for each kind of task. In this paper, the proposed MoCLE is built on InstructBLIP as the base model.
Figure 2: Overall pipeline of MoCLE. The left part is to cluster training data (i.e., a collection of vision-language tasks with diverse instructions) into several groups based on their instructions via $k$-means clustering algorithm. On the right is the MoCLE framework. $\mathbf{x}$ are the hidden representations of the sample (e.g., a sequence of tokens for images, instructions and the raw text inputs in the LLM). The input $\mathbf{x}$ is handled by 3 modules, including the universal expert, the task expert with the largest gating value, and the linear module of the LLM. After that, the outputs of these modules are aggregated and passed to the next layer. Only experts, the cluster embedding and the gate parameters are updated during training, while parameters of the LLM are frozen. We insert such structure into the query and value projection of self-attention blocks of each layer by default.
Mixture of Experts.
@@
MoE models [20, 21, 47] are considered as an effective way to increasing the model capacity in terms of parameter size. In MoE, certain parts of the model are activated while the computation is kept the same or close to its dense counterpart. Recently, it has been thoroughly investigated in the field of computer vision [44, 65], natural language processing [24, 12] and multi-modal learning [41, 48]. Among those studies, MoCE [65] tackles negative transfer in the self-supervised visual pre-training [6, 7] using a mixture of cluster experts. Different from MoCE that needs to be finetuned on downstream tasks, the proposed MoCLE method studies the zero-shot generalization to unseen tasks. Another line of works combines MoE with adapters. AdaMix [54] randomly selects adapters during training and uses the average parameters of those experts during the inference. Zadouri et al. [60] proposes mixture of LoRAs for instruction tuning of LLMs, achieving comparable results with the fully-finetuned counterpart. However, neither of them explicitly addresses task conflicts and generalization to unseen tasks, which is the focus of our work.
@@ 3 Methodology
In this section, we start with the formulation of LVLM instruction tuning and the analysis of the limitations of task experts. We then introduce the proposed MoCLE in detail. Figure 2 shows the overall framework of MoCLE.
@@ 3.1 Problem Formulation
Suppose that there are a set of datasets which are divided into held-in and held-out datasets, a large vision-language model is first fine-tuned on the held-in and then evaluated in a zero-shot manner on the held-out datasets. To unify input-output formats and promotes generalization in instruction tuning, several task templates $\{T_{i}\}$ are designed to wrap the raw inputs, i.e., a pair of text and image $(X_{\text{txt}},X_{\text{img}})$ from the corresponding dataset, where $T_{i}$ denotes the $i$-th template. For example, “Given the image, answer the question with no more than three words. {Question}” is a template for the visual question answering tasks. An instruction $I\equiv T_{i}(X_{\text{txt}})$ wraps the text input using the template.
@@ 3.2 Limitation of Task Experts
As observed in Figure 1, allocating task experts based on the categories of the training datasets brings positive transfer to similar downstream tasks. However, there are several limitations of this solution. The first limitation is that taxonomies such as “vqa” and “cap” require delicate human designs with expertise in the task performed. The second limitation is that when a new task comes, we do not know which expert to use. Further, with tasks trained in an isolated manner, knowledge sharing among different tasks is prevented, compromising generalization to unseen tasks and instructions as some might benefit from multiple tasks. Hence, we propose the MoCLE in the following sections to alleviate these two limitations.
@@ 3.3 Clustering Data by Instructions
The goals of partitioning the training data are two-fold. We target to train a task expert within a collection of similar tasks so as to prevent task conflicts. Meanwhile, we expect new tasks to be assigned to the proper experts based on its cluster without manual intervention. To achieve these, we should find a proper description for the task. Instructions serve as the foundation for LVLMs because they encapsulate the commands and intentions that guide responses of the model. The instructions $I$ in LVLMs consist of templates and the raw inputs. By understanding the templates in the instructions, the model can discern the type of task it needs to perform. The raw inputs further infuse this framework with the substantive content that the model must process and respond to. Therefore, by conducting clustering on the instructions, we can partition training data without human labor into several clusters, each containing a set of similar tasks. We further verify this through experiments later in Sec. 4.5. Moreover, even when novel instruction tasks come, we can obtain a cluster index for them using the pretrained clustering model.
Formally, denoting $\mathcal{E}(\cdot)$ as a pretrained sentence encoder, $\mathbf{e}_{i}=\mathcal{E}(I_{i})$ is the sentence representation of an instruction $I_{i}$, we use the $k$-means clustering algorithm to group all the instructions in the training datasets into $K$ clusters by iteratively minimizing $\sum_{j=1}^{K}\sum_{\mathbf{e}_{i}\in S_{j}}\|\mathbf{e}_{i}-\mathbf{c}_{j}\|^{2}$, where $S_{j}$ denotes the set of instructions assigned to the $j$-th cluster and $\mathbf{c}_{j}$ denotes the centroid of the $j$-th cluster. In each iteration of the $k$-means clustering algorithm, each instruction is assigned to the nearest centroid with all centroids updated to the average of instructions in the corresponding cluster.
3.4 Mixture of Cluster-conditional Experts
In Sec. 1, we defined a static one-to-one mapping between the task categories and expert parameters. For example, the data in the $i$-th cluster are processed by expert-$j$ for any layer at all the time. Such static mapping is sub-optimal since $N$ tasks categories require $N$ task experts, resulting prohibited model sizes as $N$ becomes large. In fact, an expert does not necessarily have to handle only one kind of tasks all the time. That is, at different layers, a task could have different choices of experts. Concretely, consider $L$ layers each with $N$ experts, then there are $N^{L}$ number of choices in total. This would accommodate much more tasks while keep $N$ small. Yet, under such a circumstance, how to choose the expert for a given task remains unclear.
We propose Mixture of Cluster-conditional LoRA Experts (MoCLE) that learns to activate the LoRA expert at each layer given the cluster of the data. Specifically, denoting by $E$ the number of experts, we introduce a gate vector $\mathbf{G}\in\mathbb{R}^{E}$ that, given any input $\mathbf{x}_{i}$, determines the experts to which the input is routed. The gate vector is calculated as,

### 22 Mar 2024 v4

# Project Page: https://gyhdog99.github.io/projects/mocle/

@@ Abstract
Instruction tuning of Large Vision-language Models (LVLMs) has revolutionized the development of versatile models with zero-shot generalization across a wide range of downstream vision-language tasks. However, the diversity of training tasks of different sources and formats would lead to inevitable task conflicts, where different tasks conflict for the same set of model parameters, resulting in sub-optimal instruction-following abilities. To address that, we propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel Mixture of Experts (MoE) architecture designed to activate the task-customized model parameters based on the instruction clusters. A separate universal expert is further incorporated to improve generalization capabilities of MoCLE for novel instructions. Extensive experiments on 11 zero-shot tasks demonstrate the effectiveness of MoCLE.
Keywords:
Instruction tuning Mixture of Experts Vision and language
@@ 1 Introduction
There has been a continuously increasing trend to develop intelligent assistants that can follow human instructions [4, 48, 8], with instruction tuning emerging as a notably effective approach. This method leverages large-scale well-formatted instruction data to empower Large Language Models (LLMs) to execute various human instructions, showcasing their ability to generalize across novel unseen tasks [41]. Likewise, efforts have been made to introduce similar capabilities to Large Vision-language Models (LVLMs) [3, 66, 63, 10, 5], including LLaVA series [38, 37], MiniGPT-4 [73] and InstructBLIP [14]. They connect a trainable visual encoder with a frozen LLM and endow the LVLMs with the instruction-following ability by finetuning on diverse multi-modal instruction datasets.
@@
It is observed in both the LLMs [51, 59, 13] and LVLMs [3, 69, 31] that the ability to generalize to novel unseen instructions necessitates multi-task instruction tuning, i.e., training on a diverse collection of instruction-following tasks. However, the complexity of various instruction tasks brings difficulties for the model fine-tuning. Specifically, [60] find that under certain model sizes, multi-task instruction tuning even fails to bring performance gains for zero-shot tasks compared to the original models, mainly attributed to the negative transfer phenomenon [68] during multi-task instruction tuning, where the model struggles to optimize the losses of multiple conflicted tasks, leading towards the sub-optimal experimental performance for these tasks.
@@
Figure 1: Performance of the instruction-finetuned LVLMs on zero-shot tasks, where larger values indicate better performance. Only 2 out of 7 tasks benefit from instruction tuning from all the data, while the task experts show better performance on the other 5 tasks (i.e., Flickr 30K, GQA, HM, SciQA and IconQA).
@@
Similarly, tasks for vision-language instruction tuning (e.g., visual question answering and image captioning) focus on different perspectives of LVLMs, and thus, resulting in conflicts when sharing all the parameters as most studies adopt. In our preliminary study, we first split the instruction data into two disjoint subsets, including “cap” for image captioning and “vqa” for visual question answering. We then train InstructBLIP [14] using LoRA [21] on 3 kinds of data: “cap”, “vqa” and “full” (the full data) separately, to obtain 3 set of parameters (i.e., task experts). Following the held-out evaluation protocol [14], we evaluate these experts on the unseen datasets/tasks with the best expert. As shown in Figure 1, on 5 out of the 7 downstream tasks, the InstructBLIP instruction-tuned on all the data is outperformed by the task expert finetuned with only a subset of data. Among the 5 tasks, the Flickr30k belongs to “cap”, and SciQA, GQA and IconQA belong to “vqa”, revealing that instruction tuning on the similar tasks brings positive transfer to downstream tasks, while training on the full data with dissimilar tasks could hurt the generalization performance.
@@
The disjoint task expert above is a naïve solution to negative transfer, where we manually partition training tasks and train each expert individually. However, there are several limitations of this strategy: (1) The taxonomies such as “vqa” and “cap” require human expertise and become difficult to scale as the number of tasks grows. (2) The ability to generalize to unseen tasks is inhibited, as we do not know which expert to choose for novel tasks, while some new tasks might benefit from multiple training tasks (e.g., VSR and TextVQA as in Figure 1). In this regard, specialization and generalization of LVLMs becomes a dilemma.
@@
Our work aims to develop an automatic and practical partition strategy and a network architecture that can strike a balance between specialization and generalization. We propose Mixture of Cluster-conditional LoRA Experts (MoCLE) for vision-language instruction tuning. Under the proposed MoCLE framework, we first cluster instructions of all the training data into several centroids via a pre-trained clustering model. In this way, similar tasks that could bring positive transfer to each other are automatically grouped into the same cluster, and different tasks that may cause conflict are separated (more justification for the usage of instruction clusters is detailed in Sec. 3.3). Then we construct several task experts with each concentrating on the specific cluster. Using the cluster as a condition, a router dispatches the input data to one of the specialized task experts and a universal expert that is shared among all the data. As we activate a specialized expert for a group of similar tasks, tasks that are less alike are learned via separate experts, mitigating task conflicts. Meanwhile, since the universal expert trained on all tasks also contributes to the model outputs, we can enjoy generalization and specialization simultaneously.
@@
We validate effectiveness of MoCLE on the state-of-the-art InstructBLIP [14]. MoCLE achieves significant improvement on the held-out tasks [14] including the image captioning (e.g., Flickr30K [64]), the text-based visual question answering (e.g., TextVQA [56]), the knowledge-based visual question answering (e.g., ScienceQA [42]), and novel visual reasoning (e.g., Visual Spatial Reasoning [36]).
@@
    We identify the negative transfer phenomenon[40, 71] of task conflicts during the instruction tuning of LVLMs.
@@
    We propose Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel parameter-efficient finetuning framework suitable for the vision-language instruction tuning, to mitigate the task conflicts and enjoy the benefits of huge data training simultaneously.
@@
    Our proposed MoCLE achieves remarkable performance gains on the unseen held-out tasks compared with a strong baseline model InstructBLIP [14].
@@ 2 Related Work
2.0.1 Multi-task Instruction Tuning.
@@
Instruction tuning, as elucidated in [51, 60], is a technique where a language model is finetuned over a collection of tasks and instruction templates that tells the model the intention of the task. The primary goal of this technique is to teach the model to understand the relationship between the instruction and the input/output pairs, so that the model could generalize to unseen task given novel instructions. Increasing the number of diverse instructions [51], the number of tasks [59, 13], and diversity of data [72] have been shown to improve the performance. However, [60] observe under certain model sizes, instruction tuning of language models fail to bring performance gains to unseen tasks when compared with untuned base models. They argue this is due to the fact all model capacity is used to learn the mixture of instruction tuning tasks. In this work, we aim to mitigate such issue in vision-language instruction tuning by the specialized experts.
2.0.2 Large Multi-modal Models.
@@
To inherit strong reasoning and generalization abilities of LLMs and meanwhile understand visual inputs, recent works including Flamingo [2], BLIP-2 [28], LLaMA-adapter [67] connect LLMs to the visual modality with a frozen visual encoder and a trainable module, which is then finetuned with plain image-text pairs. To further improve instruction following abilities and response quality, LLaVA series [38, 37] finetune the whole LLM on the vision-language conversational data generated by GPT-4. MiniGPT-4 [73] performs training on ChatGPT-generated image captions instead of traditional caption data that are shorter and informational sparser. InstructBLIP [14] collects a wide range of academic VQA datasets and manually crafts 10 to 15 distinct instruction templates for each kind of task. In this paper, we primarily consider InstructBLIP as the base model.
2.0.3 Mixture of Experts.
@@
MoE models [23, 24, 53] are considered as an effective way to increase the model capacity in terms of parameter size. Within MoEs, certain parts of the model are activated while the computation is kept the same or close to its dense counterpart. Recently, it has been thoroughly investigated in the field of computer vision[50, 71], natural language processing[27, 15] and multi-modal learning[47, 54, 12, 11, 34]. MoCE [71] tackles negative transfer in the self-supervised visual pre-training[6, 7] using a mixture of cluster experts. Different from MoCE requiring to be finetuned on downstream tasks, our MoCLE studies the zero-shot generalization to unseen tasks. Another line of works combines the MoE with adapters. AdaMix[58] randomly selects adapters during training and adopts the average of expert parameters during inference. Zadouri et al. [65] propose mixture of LoRAs for instruction tuning of LLMs.
Concurrently, MoE-LLaVA [34] sparsifies LVLMs via MoEs for more efficient model scaling. The most relevant to ours include the Octavius [12] and LLaVA-MoLE [11], which also address task conflicts in vision-language instruction tuning by applying the MoE architecture with LoRAs. However, they do not explore the issue of instruction generalization in the architecture designs. Instead, we design a novel Mixture of Cluster-conditional LoRA Experts together with the universal experts, which is the first work addressing both task conflicts and generalization with a unified architecture.
@@ 3 Methodology
In this section, we start with the formulation of LVLM instruction tuning and the analysis of the limitations of task experts. We then introduce the proposed MoCLE in detail, with the overall framework shown in Figure 2.
Figure 2: Overall pipeline of MoCLE. The left part is to cluster training data (a collection of vision-language tasks with diverse instructions) into several groups based on the instructions via $k$-means clustering algorithm. On the right is the MoCLE framework. $\mathbf{x}$ is the hidden representation of the sample (e.g., a sequence of tokens for images, instructions and the raw text inputs in the LLM) which is handled by 3 modules, including the universal expert, task expert, and the original linear module of the LLM. After that, the outputs of these modules are aggregated and passed to the next layer. Only experts, the cluster embedding and the gate parameters are updated during training, while parameters of the LLM are frozen. We insert such structure into the query and value projection in self-attention blocks of each layer by default.
@@ 3.1 Problem Formulation
Suppose that there is a set of datasets that are divided into held-in and held-out datasets [14], where a large vision-language model is first finetuned on the held-in and then evaluated on the held-out datasets in a zero-shot manner. To unify and diversify input-output formats and promotes in instruction tuning, several task templates $\{T_{i}\}$ are designed to wrap the raw inputs, a pair of text and image $(X_{\text{txt}},X_{\text{img}})$ from the corresponding dataset, where $T_{i}$ denotes the $i$-th template. For example, “Given the image, answer the question with no more than three words. {Question}” is a template for visual question answering tasks. The instruction is defined as $I\equiv T_{i}(X_{\text{txt}})$ that wraps text inputs using the template.
@@ 3.2 Limitation of Task Experts
Figure 1 indicates that allocating task experts based on the categories of training datasets brings positive transfer to similar downstream tasks. Nevertheless, this approach has several notable limitations. First, taxonomies such as “vqa” and “cap” require delicate human designs with expertise in the task performed. Another limitation is the challenge of selecting the appropriate expert for a newly introduced task when testing. Further, with tasks trained in an isolated manner, knowledge sharing among different tasks is prevented, compromising generalization to unseen tasks and instructions as some might benefit from multiple tasks. Hence, we propose MoCLE in the following sections to address these limitations.
@@ 3.3 Clustering Data by Instructions
The purposes of partitioning the training data are two-fold. We target at training a task expert within a collection of similar tasks so as to prevent task conflicts. Meanwhile, we expect novel tasks to be assigned to the proper experts based on their cluster without manual intervention. To achieve these, we should find a proper description for the task. Instructions serve as the foundation for LVLMs because they encapsulate the commands and intentions that guide responses of the model. The instructions $I$ in LVLMs consist of templates and the raw inputs. By understanding the templates in the instructions, the model can discern the type of task it needs to perform. The raw inputs further infuse this framework with the substantive content that the model must process and respond to. Therefore, by conducting clustering on the instructions, we can partition training data without human labor into several clusters, each containing a set of similar tasks, which is further verified through experiments in Sec. 4.5. Moreover, even when novel instruction tasks come, we can obtain a cluster index for them using the pre-trained clustering model without human intervention.
Formally, denoting $\mathcal{E}(\cdot)$ as a pre-trained sentence encoder, $\mathbf{e}_{i}=\mathcal{E}(I_{i})$ is the sentence representation of an instruction $I_{i}$, we use the $k$-means clustering algorithm to group all instructions in the training datasets into $K$ clusters by iteratively minimizing $\sum_{j=1}^{K}\sum_{\mathbf{e}_{i}\in S_{j}}\|\mathbf{e}_{i}-\mathbf{c}_{j}\|^{2}$, where $S_{j}$ denotes the set of instructions assigned to the $j$-th cluster and $\mathbf{c}_{j}$ denotes the centroid of the $j$-th cluster. In each iteration of the $k$-means clustering algorithm, each instruction is assigned to the nearest centroid with all centroids updated as the average of the instruction representations in the corresponding cluster.
3.4 Mixture of Cluster-conditional LoRA Experts
In addition to considerations at the data level, we also suggest an architectural design to tackle the issue of negative transfer. We propose Mixture of Cluster-conditional LoRA Experts (MoCLE) that learns to activate the LoRA expert at each layer given the cluster of the data. Specifically, denoting $E$ as the number of experts, we introduce a gate vector $\mathbf{G}\in\mathbb{R}^{E}$ that, given any input $\mathbf{x}_{i}$, determines the experts to which the input is routed. The gate vector is calculated as,